OpenStudy (anonymous):

What Is The Solution To The System? 3x+y+z=7 x+3y-z=13 y=2x-1 A.)(7,13,-1) B.)(3,2,-2) C.)(2,3,-2) D.)(-2,3,2)

OpenStudy (anonymous):

plug equation 3 into 1 and 2 : 3x + (2x-1) + z = 7 x + 3(2x-1) -z = 13 5x + z = 8 7x - z = 16 add the equations : 12x = 24 x=2 plug x=2 into y=2x-1 to get: y = 3 plug x=2 into either 5x + z = 8 or 7x - z = 16 to get: 14 -z = 16 z = -2 (2,3,-2)

OpenStudy (anonymous):

Thanks